The sum of two numbers is 42. The larger number is 8 more than the smaller number. What are the two numbers?
Answer by ikleyn(52914)

.
The sum of two numbers is 42. The larger number is 8 more than the smaller number. What are the two numbers?
----------------------------------------------

Let  x  be a smaller number. 
Then the greater number is  (x+8),  according to the condition.

Their sum is  42,  which gives you an equation

x + (x+8) = 42,   or

2x + 8 = 42,   or

2x = 42 - 8 = 34.

Hence,  x = 17. 

Thus the smaller number is  17.  The greater number is 17+8 = 25.

Please check the solution yourself.